OpenStudy (anonymous):

Tina had $47.10 in her bank account on Monday. She withdrew $25 on Monday and also made a deposit of an amount d that she cannot remember. Tina had $97.43 in her account on Tuesday. Which equation can be used to solve for d, the amount she deposited on Monday?

OpenStudy (anonymous):

A. 47.10 – 25 + d = 97.43 B. 47.10 – 25 – d = 97.43 C. 47.10 + 25 – d = 97.43 D. 47.10 + 25 + d = 97.43

OpenStudy (anonymous):

This is my last question in THIS assessment...

OpenStudy (anonymous):

It would be A. because she has 47.50 subtracts/withdraws 25 and adds d the unknown amount and has a total of 97.43 on Tuesday.

OpenStudy (anonymous):

*47.10

OpenStudy (anonymous):

Thank you @xxyourenotalonexx by the way, I really like your username.

OpenStudy (anonymous):

No problem and thank you :D it's a line from one of my favourite songs.

OpenStudy (anonymous):

Oh! What song?

OpenStudy (anonymous):

I got 100% On my test C:

OpenStudy (anonymous):

It's You're Not Alone by Of Mice & Men

OpenStudy (anonymous):

Also, congratulations!

OpenStudy (anonymous):

Okey, I'll listen to it!!

OpenStudy (anonymous):

It's screamo… just warning you. xD

OpenStudy (anonymous):

And could you help me out with my question?